# Symbolic Constraints in Polyhedral Enclosure and Tetrahedral Decomposition in Genus-0 Polyhedra

### v6

@@ 4.2 External Angle Unit Bounds
$M_{\max}=10(V-2),\qquad M_{\min}=\begin{cases}4V+6,&\text{$V$ odd},\\ 4V+4,&\text{$V$ even}.\end{cases}$
4.3 Configuration Series over Flatness
From Proposition 4.1, each unit decrease in the flatness measure $S$ increases both $F$ and $E$ by $1$. Starting from the extremal values in Corollary 4.2, I have:
$S_{k}=S_{\max}-k,\quad F_{k}=F_{\min}+k,\quad E_{k}=E_{\min}+k,$
for $k=0,1,\dots,S_{\max}$.
@@ 4.4 Counting Distinct Combinatorial Configurations
$\text{Combinations}=S_{\max}+1=\begin{cases}\frac{3}{2}V-\tfrac{11}{2},&\text{$V$ odd},\\ \frac{3}{2}V-5,&\text{$V$ even}.\end{cases}$
Heuristic 4.1 (External bounds and parity effects).
The expressions for the extremal ranges in §3 (including $F_{\min}$, $E_{\min}$, $S_{\max}$ and the parity-sensitive offsets) follow from Proposition 4.1 together with Lemma 4.1 and Corollary 4.1. They are supported by exhaustive checks for small $V$ and by constructive families achieving the bounds, but are not proved to be tight in full generality.
@@ 4.5 Triangulation Sequences and the Invariance of Euler’s Characteristic
Proposition 4.3 (Triangulated extrema and invariance of $\chi$).
For a genus-zero polyhedral surface composed entirely of triangles,
$F_{\max}=2(V-2),\qquad E_{\max}=3(V-2).$
$F(S)=F_{\max}-S,\qquad E(S)=E_{\max}-S,\qquad V-E(S)+F(S)=2.$
Proof.
Each triangle has three edges and each edge is shared by two triangles, so $2E_{\max}=3F_{\max}$. Substituting $E_{\max}=\tfrac{3}{2}F_{\max}$ into Euler’s formula $V-E_{\max}+F_{\max}=2$ yields $F_{\max}=2(V-2)$ and $E_{\max}=3(V-2)$. A coplanarity merge of two adjacent triangles removes one surface edge and one triangle, so $F\mapsto F-1$ and $E\mapsto E-1$ with $V$ unchanged. Hence VE+FV-E+F is invariant and equal to $2$ for all admissible $S$. ∎
@@ 5 Internal Decomposition and Tetrahedral Estimation
Section 5 shifts to internal volumetric decomposition and presents a clean identity linking tetrahedra, gluing planes, and internal segments. It first explores the internal volumetric structure of valid polyhedra, defined by a given vertex count $V$ and the surface properties derived in previous sections. Specifically, it aims to estimate the number or range of internal tetrahedra ($T$) that can fill a polyhedron, based on its triangulated faces and edge configurations. This continues the logic of the preceding analysis, shifting from surface validation to interior decomposition. The empirical formulas presented here result from the strict application of the tetrahedral decomposition process outlined in the Methods section.
@@
Throughout this section, “SALT+MIE” denotes a restricted construction class (genus–0, normal form; shell layers; no edge bisections). Statements labeled as identities are exact under the stated hypotheses; ranges and patterns specific to SALT+MIE are not claimed to hold for arbitrary tetrahedralizations.
5.1 Internal Structure Variables
Let a polyhedron with $V\geq 4$ vertices be internally composed of:
  • $T$: Number of tetrahedra,
  • $S_{i}$: Number of internal non-intersecting triangulation segments to fully triangulate internal gluing surfaces,
  • $N_{i}$: Number of internal gluing triangle units.
5.2 Lower bounds and restricted upper bounds (SALT+MIE)
Under SALT with the Minimal Interior Edge rule (at most one new interior edge per layer), the exact identities of Propositions 6.26.3 imply
$T_{\min}=V-3\quad(S_{i}=0),\qquad T=V-3+S_{i},\qquad N_{i}=2T-V+2=V-4+2S_{i}.$
Because at most one interior edge is introduced per layer and the first and last layers cannot host a new interior edge, one has $0\leq S_{i}\leq V-5$, whence
$T_{\max}=V-3+(V-5)=2(V-4),\qquad S_{i,\max}=V-5,\qquad N_{i,\max}=V-4+2(V-5)=3(V-4)-2.$
5.3 SALT ladder family
Within SALT+MIE, a transparent subfamily is the “ladder” construction where successive nontriangular boundary faces are filled one by one. For $k=0,\ldots,V-5$,
$T_{k}=V-3+k,\quad N_{i,k}=V-4+2k,\quad S_{i,k}=k,$
and $T_{k}-N_{i,k}+S_{i,k}=1$ holds exactly.
5.4 Count of Valid Internal Configurations
$\text{Count of valid }(T,N_{i},S_{i})\text{combinations}=V-4$
Heuristic 5.1 (Internal ranges and configuration ladder).
The ranges in §4 for $(T_{\min},T_{\max})$, $(N_{i,\min},N_{i,\max})$, and $S_{i,\max}$, together with the ladder
Empirical surface-type filtering at fixed $V$ (small-$V$ compression).
The ladder identities determine the full algebraic family $T=V-3+S_{i}$ and $N_{i}=V-4+2S_{i}$ under SALT+MIE. However, the realizability of the upper rungs depends on the external surface type $(E,F,S)$ at fixed $V$. Table LABEL:tab:V8-filter (compiled for $V=8$) exhibits a sharp compression: as external flatness $S$ increases (equivalently, as $E$ decreases so that E+S=3V6E+S=3V-6 remains fixed), progressively fewer $(T,N_{i},S_{i})$ rungs are realized.
In particular, for $S=6$ (the cube), only the two configurations with the lowest tetrahedral counts occur: $T=5$ (the $S_{i}=0$ rung) and $T=6$ (the $S_{i}=1$ rung). The higher algebraically admissible rungs $T=7$ and $T=8$ do not appear, despite being allowed by the ladder identities. Thus, the observed maximum is $T=6$, not the algebraic ceiling $T=8$.
This suggests that at large $S$, internal gluing structure may vary substantially in planarity, allowing non-coplanar interna gluingl triangles $N_{i}$ while maintaining minimal internal segmentation $S_{i}$, but that this flexibility is strongly constrained in small polyhedra. For small $V$, the interaction between surface structure and layering constraints therefore appears to impose a much sharper upper bound on $T$ than is predicted by the global incidence identities alone.
5.5 Example for $V=8$
$E$ $S$ (External) $T$ (Tetrahedra) $N_{i}$ (Triangles) $S_{i}$ (Segments)
18 0 5 4 0
18 0 6 6 1
18 0 7 8 2
18 0 8 10 3
17 1 5 4 0
17 1 6 6 1
17 1 7 8 2
16 2 5 4 0
16 2 6 6 1
15 3 5 4 0
15 3 6 6 1
14 4 5 4 0
14 4 6 6 1
13 5 5 4 0
13 5 6 6 1
12 6 5 4 0
12 6 6 6 1
Note. Here $E$ denotes the number of boundary edges prior to triangulating non-triangular faces. For fixed $V=8$, the triangulated boundary satisfies E+S=3V6=18E+S=3V-6=18.
  • When $S=0$: all V4V-4 ladder configurations are realized.
  • When $S=1$: the top ladder rung $(S_{i}=S_{i,\max})$ is not realized.
Surface flatness $S$ thus correlates with increased constraint on internal decomposition at small $V$, but it does not act as an independent control parameter. Instead, realizability appears to be governed by the necessary packing of internal tetrahedra under fixed $(V,N)$ constraints. For example, even with a fully triangulated surface ($S=6$), the cube admits only $T=5$ or $T=6$ tetrahedra under SALT+MIE.
These behaviors are likely to become increasingly complex as $V$ increases, due to the geometric and combinatorial expansion of allowable internal connectivity. Accordingly, while empirical tendencies can be identified at fixed small $V$ (e.g., $4\leq V\leq 8$), extrapolating them to larger polyhedra is likely to be misleading, as it obscures emergent phenomena such as multilayer volumetric structure.
@@ 5.6 Determining admissible tetrahedral ranges from polygonal surface data
We consider the following problem: given a closed genus–0 polyhedral surface specified by a multiset of polygonal faces, determine the admissible range of tetrahedral decompositions consistent with purely symbolic (incidence-based) constraints. Throughout, we restrict attention to normal–form decompositions: all vertices lie on the boundary, no interior vertices are introduced, and all tetrahedra are glued face-to-face. When we invoke SALT+MIE, we mean the further restricted construction class used in §4 (shell-aligned layering; no edge bisections; at most one new interior segment per shell layer).
@@ 5.6.1 Input data and boundary normalization
which counts the number of triangulation segments required to triangulate the surface without introducing new vertices.
After triangulation, the boundary data become
@@ 5.6.4 Solution family and feasibility constraints
Equations (1) and (3) parameterize all admissible triples $(T,N_{i},S_{i})$ by the single integer $T$. Solving (1) for $N_{i}$ and substituting into (3) yields
@@ 5.6.5 Restricted SALT+MIE ranges (construction-based)
Remark (Toward realizability criteria beyond the SALT ladder).
The results above distinguish three layers of constraint on internal tetrahedralization: exact incidence identities, construction-based restrictions (SALT+MIE), and empirical realizability observed in small cases. The first two layers generate a complete algebraic ladder of admissible $(T,N_{i},S_{i})$ configurations at fixed $V$, but they do not determine which rungs are realized for a given surface type $(E,F,S)$.
The $V=8$ examples show that this gap is substantive rather than technical. Although all ladder rungs are permitted algebraically, realizability depends strongly on how surface flatness is distributed across the boundary. For small polyhedra this produces sharp truncations of the ladder, while for larger $V$ similar constraints may degrade gradually, giving rise to multilayer interior structure rather than hard cutoffs. This is yet to be explored.
6 Inferring External Structure from Internal Decomposition and Euler’s Characteristic
This section relates the internal volumetric decomposition of a genus-0 polyhedron into tetrahedra to the combinatorics of its boundary surface. We first record an exact topological identity that holds for any tetrahedralization of a 33–ball, then specialize it to the normal–form setting used throughout this paper. We then present heuristic incidence formulas that allow one to infer $V$, $E$, and $F$ from $(T,N_{i},S,S_{i})$, and an “extended Euler-type” identity that follows under those heuristic assumptions.
6.1 Exact topological Euler link
@@
Let $T$ be the number of tetrahedra in a genus-0 tetrahedralization, $N_{i}$ the number of internal gluing triangles, $E_{i}$ the number of interior edges, and $V_{i}$ the number of interior vertices. Then
$\boxed{T-N_{i}+E_{i}-V_{i}=1}$
always holds.
Proof.
Count cells in the tetrahedral complex:
  • Vertices: $V+V_{i}$ (boundary plus interior),
  • Edges: $E_{\triangle}+E_{i}$, where $E_{\triangle}$ are the edges of the triangulated boundary,
  • Faces: $N+N_{i}$, where $N$ is the number of boundary triangular faces (angle units),
  • 3–cells: $T$ tetrahedra.
Euler’s formula for a 33–ball gives
$(V+V_{i})-(E_{\triangle}+E_{i})+(N+N_{i})-T=1.$
Boundary Euler ($V-E_{\triangle}+F_{\triangle}=2$) with $F_{\triangle}=N$ eliminates the boundary terms, yielding
$2-E_{i}+N_{i}-T+V_{i}=1,$
which rearranges to $T-N_{i}+E_{i}-V_{i}=1$. ∎
Normal-form specialization.
In the normal form adopted here:
  • All vertices lie on the boundary $\Rightarrow V_{i}\equiv 0$,
  • Every interior edge is an internal triangulation segment $\Rightarrow E_{i}=S_{i}$.
Thus the exact identity reduces to
$\boxed{T-N_{i}+S_{i}=1}$
for all constructions considered in this manuscript.
@@
Let a genus–0 polyhedron in normal form be tetrahedralized into $T$ tetrahedra with $N_{i}$ internal gluing triangles. After triangulating all boundary faces, the number of boundary triangles (angle units) is $N=F_{\triangle}=F+S=2V-4$. Counting triangle units gives
$4T-2N_{i}=N=2V-4\quad\Longrightarrow\quad\boxed{\,N_{i}=2T-V+2\,}.$
Proposition 6.3 (Exact 33D Euler link; normal-form specialization).
@@
For any tetrahedral complex filling a 33–ball, $T-N_{i}+E_{i}-V_{i}=1$. In normal form $V_{i}=0$, and every interior edge is an interior triangulation segment; define $S_{i}:=E_{i}$. Then

### v7

@@ 4.3 External Angle Unit Bounds
$M_{\mathrm{ub}}=10(V-2),\qquad M_{\mathrm{lb}}=\begin{cases}4V+6,&\text{$V$ odd},\\ 4V+4,&\text{$V$ even}.\end{cases}$
@@ 4.4 Counting Distinct Combinatorial Configurations
$\text{Combinations}=S_{\mathrm{ub}}+1=\begin{cases}\frac{3}{2}V-\tfrac{11}{2},&\text{$V$ odd},\\ \frac{3}{2}V-5,&\text{$V$ even}.\end{cases}$
Heuristic 4.1 (Realizability across intermediate flatness).
Proposition 4.1 and Lemma 4.1 imply the admissible parity-sensitive ranges for $E$, $F$, and $S$ in Corollary 4.2. What remains heuristic in this work is the claim that every integer $S$ in the admissible range is realizable by a simple 33–connected planar graph in normal form; see Heuristic Proposition 4.3.
@@ 5 Internal Decomposition and Tetrahedral Estimation
This section shifts from surface validation to internal volumetric structure. We record exact incidence identities governing tetrahedral decompositions of genus–0 polyhedra in normal form, then study how restricted construction rules (SALT+MIE) and surface structure constrain realizability.
@@
Throughout, we restrict attention to normal–form decompositions: all vertices lie on the boundary, no interior vertices are introduced, and all tetrahedra are glued face-to-face. When we invoke SALT+MIE, we mean the further restricted construction class used in §4 (shell-aligned layering; no edge bisections; at most one new interior segment per shell layer).
@@ 5.1 Determining admissible tetrahedral ranges from polygonal surface data
We consider the following problem: given a closed genus–0 polyhedral surface specified by a multiset of polygonal faces, determine the admissible range of tetrahedral decompositions consistent with purely symbolic (incidence-based) constraints.
@@ 5.1.1 Input data and boundary normalization
which equals the number of diagonals in any triangulation of the polygonal faces (without introducing new vertices).
Since triangulating a face of degree $k$ introduces exactly k3k-3 diagonals, summing over faces gives $S$. After triangulation, the boundary data become
@@ 5.1.3 Exact incidence identities (normal form)
Appendix B demonstrates that the identity (3) is robust when internal structure arises solely from boundary-induced triangulation (the SALT+MIE setting), and that it fails once interior Steiner points are introduced. Appendix C, by contrast, demonstrates invariance under surface-embedded edge Steiner points.
@@ 5.1.4 Solution family and feasibility constraints
Equations (1) and (3) parameterize all incidence-admissible triples $(T,N_{i},S_{i})$ by the single integer $T$. Solving (1) for $N_{i}$ and substituting into (3) yields
@@ 5.1.5 Restricted SALT+MIE ranges (construction-based)
5.2 Exact interior incidence identities (normal form)
@@
Let $T$ be the number of tetrahedra in a genus–0 tetrahedralization, $N_{i}$ the number of internal gluing triangles, $E_{i}$ the number of interior edges, and $V_{i}$ the number of interior vertices. Then
$\boxed{T-N_{i}+E_{i}-V_{i}=1.}$
@@
Let a genus–0 polyhedron in normal form be tetrahedralized into $T$ tetrahedra with $N_{i}$ internal gluing triangles. After triangulating all boundary faces, the number of boundary triangles is N=2V4N=2V-4. Counting triangle units gives
$4T-2N_{i}=N\quad\Longrightarrow\quad\boxed{N_{i}=2T-V+2.}$
Proposition 5.3 (Normal-form specialization).
@@
In normal form, $V_{i}\equiv 0$ and every interior edge is an internal triangulation segment; define $S_{i}:=E_{i}$. Then
$\boxed{T-N_{i}+S_{i}=1,}$
and combining with Proposition 5.2 yields
$\boxed{T=V-3+S_{i}.}$
These identities are exact and independent of any construction method.
5.3 General solution family and incidence bounds
For fixed boundary data (hence fixed $V$), Propositions 5.25.3 parameterize all admissible triples $(T,N_{i},S_{i})$ by the integer variable $T$, subject only to nonnegativity. In particular,
$N_{i}(T)=2T-V+2,\qquad S_{i}(T)=T-V+3.$
From $N_{i}\geq 0$ one obtains the general incidence bound
$T\geq\left\lceil\frac{V-2}{2}\right\rceil,$
which holds for any normal-form tetrahedralization.
5.4 SALT ladder family
Within SALT+MIE, a transparent subfamily is the “ladder” construction where a polygon is decomposed into tetrahedra by the identification of internal triangular gluing faces one by one. For $k=0,\ldots,V-5$,
$T_{k}=V-3+k,\quad N_{i,k}=V-4+2k,\quad S_{i,k}=k,$
and $T_{k}-N_{i,k}+S_{i,k}=1$ holds exactly.
For fixed $V$, the ladder construction realizes exactly V4V-4 distinct admissible triples $(T,N_{i},S_{i})$ under SALT+MIE.
Heuristic 5.1 (Internal ranges and configuration ladder).
The ranges for $(T_{\min},T_{\max})$, $(N_{i,\min},N_{i,\max})$, and $S_{i,\max}$, together with the ladder
5.5 Empirical truncation at fixed $V$
Although the ladder exhausts the algebraic SALT+MIE range, realizability at fixed $V$ depends on the external surface structure $(E,F,S)$. Table 2 (for $V=8$) exhibits a sharp compression: as surface flatness $S$ increases, progressively fewer ladder rungs are realized. For example, when $S=6$ (the cube), only $T=5$ and $T=6$ occur, despite the algebraic allowance of $T=7,8$. Thus the effective maximum is strictly below the incidence ceiling.
Table 2: Valid $(T,N_{i},S_{i})$ configurations observed at $V=8$ under SALT+MIE, grouped by external flatness $S$. Here $E$ denotes the number of boundary edges prior to triangulating non-triangular faces. For fixed $V=8$, the triangulated boundary satisfies E+S=3V6=18E+S=3V-6=18.
$E$ $S$ $T$ $N_{i}$ $S_{i}$
18 0 5 4 0
18 0 6 6 1
18 0 7 8 2
18 0 8 10 3
17 1 5 4 0
17 1 6 6 1
17 1 7 8 2
16 2 5 4 0
16 2 6 6 1
15 3 5 4 0
15 3 6 6 1
14 4 5 4 0
14 4 6 6 1
13 5 5 4 0
13 5 6 6 1
12 6 5 4 0
12 6 6 6 1
  • When $S=0$: all V4V-4 ladder configurations are realized.
  • When $S=1$: the top ladder rung $(S_{i}=S_{i,\max})$ is not realized.
This compression is empirical and specific to $V$; it reflects the interaction between boundary structure and volumetric packing rather than a failure of the incidence identities.
5.6 Flexibility and Inflexibility of Internal Arrangements at Fixed $V$
The preceding sections establish admissible ranges for the number of tetrahedra $T$ and internal gluing triangles $N_{i}$ for a polyhedron with fixed boundary vertex count $V$ under the SALT+MIE construction rules. However, these bounds alone do not distinguish between configurations that admit multiple internal realizations and those that are structurally rigid. In this section, we introduce a qualitative distinction between inflexible and flexible internal arrangements, based on whether the surface triangulation budget $S$ can vary while $V$ and $T$ are held fixed.
5.6.1 Definition: $S$-flexibility
Fix a boundary vertex count $V$ and a tetrahedral count $T$. We say that a configuration is $S$-inflexible if all valid SALT+MIE realizations with these fixed values of $(V,T)$ admit the same value of $S$. Conversely, a configuration is $S$-flexible if there exist two valid realizations with the same $(V,T)$ but different values of $S$.
Intuitively, $S$-inflexibility corresponds to a boundary and interior structure that forces a unique surface triangulation compatible with a valid tetrahedral fill, while $S$-flexibility indicates the presence of discrete triangulation choices that do not alter the global counts $V$ or $T$.
5.6.2 Inflexible class: bipyramidal configurations
The canonical examples of $S$-inflexible arrangements are bipyramidal configurations. These consist of two apex vertices (one “above” and one “below”) connected to all vertices of a central polygonal cycle. All faces are triangular, and the interior tetrahedra fan out from the central polygon in a forced manner.
For such configurations:
  • the induced tetrahedral adjacency structure is fixed,
  • no local rearrangement of internal gluing surfaces is possible without changing either $V$ or $T$,
  • the surface triangulation budget $S$ is uniquely determined (typically $S=0$).
An explicit example is the case $(V,T)=(8,8)$, where the only realizable configuration is a hexagonal bipyramid. In this case, no alternative triangulation of the surface exists that is compatible with a valid tetrahedral decomposition, and the configuration is strictly $S$-inflexible.
5.6.3 First deviation: single-vertex departure from bipyramidity
Flexibility first appears when the bipyramidal structure is minimally perturbed. Specifically, if one vertex leaves the equatorial polygon and is reassigned to either the upper or lower apex side, the central polygon is reduced by one vertex. One side of the polyhedron remains a pyramid, while the opposite side now contains a single quadrilateral face.
This quadrilateral face admits exactly two triangulations, yielding two distinct but valid surface configurations:
  • one realization with $S=0$,
  • one realization with $S=1$.
Crucially, both realizations share the same values of $(V,T)$. Thus, this class represents a minimally flexible arrangement, with exactly one discrete degree of freedom in $S$. The interior accommodates one fewer tetrahedron than the corresponding pure bipyramid, but remains otherwise tightly constrained.
5.6.4 Higher deviations: increasing flexibility
Additional flexibility arises when further vertices leave the equatorial polygon. For example:
  • two vertices may be reassigned to the same side, producing a pentagonal face;
  • or one vertex may move upward while another moves downward, producing a quadrilateral face on each side.
Each non-triangular face introduces an independent triangulation choice. As a result, the set of admissible values of $S$ grows, while $V$ and $T$ remain fixed. These configurations are genuinely $S$-flexible, exhibiting multiple realizable internal arrangements and a nontrivial range of surface triangulation budgets.
5.7 Interpretation
This progression suggests a natural stratification of configurations at fixed $V$:
bipyramidal (rigid) $\rightarrow\;$ single-defect (minimally flexible) $\rightarrow\;$ multi-defect (flexible).
In this view, bipyramidal configurations form the rigid core of the admissible space, while flexibility emerges precisely as vertices are removed from the equatorial cycle. Each such departure introduces a higher-degree face and hence an additional discrete triangulation choice , reflected in the admissible range of surface flatness $S$.
This distinction clarifies why the admissible SALT+MIE bounds alone do not determine $T$ or $S$ uniquely: realizability within those bounds depends sensitively on the internal combinatorial structure, not solely on boundary counts.
Remark (Incidence-Gated Transitions, Bipyramidal Rigidity, and Boundary Flexibility).
Although the analysis throughout this work is entirely symbolic, it is often helpful to imagine vertices as concrete points in space and tetrahedra as occupying volume. This geometric picture is not essential, but it helps explain why certain discrete effects appear abruptly rather than gradually.
A key point is that the equatorial “belt” in a bipyramidal configuration need not be coplanar. Coplanarity is a red herring: what actually governs admissibility is the pattern of external boundary edges and internal gluing segments. As long as these incidence relations are preserved, the belt vertices can ripple or kink freely in space without any effect on the tetrahedral count or on other boundary quantities (edges and faces). Geometry supplies continuous freedom within each combinatorial state.
From this perspective, bipyramidal configurations occupy a special extremal position. Their rigidity is geometrically obvious: two apices constrain a belt of vertices so tightly that there is essentially no slack. However, this rigidity is largely invisible at the level of surface incidence counts alone. In informal terms, bipyramids are often mistaken for “flexible” objects combinatorially, even though they are maximally constrained volumetrically. This is why, if one insists on geometric metaphors for rigidity, the usual description of inflexible objects as “squares” is misleading: squares are flexible in triangulated settings and signal at least a degree of boundary flexibility, whereas bipyramids are not.
Departures from bipyramidity occur through discrete reclassifications of vertex roles rather than smooth geometric limits. The primitive defect is not an apex becoming coplanar with the belt, but the reassignment of a belt vertex. When a single belt vertex ceases to attach symmetrically to both apices, remaining connected to only one apex while acquiring additional lateral connections, the belt shortens by one, an internal gluing segment is lost, and exactly one tetrahedron disappears. This produces the first unit of variability: two triangular faces may unite to form a boundary quadrilateral ($S=1$), the tetrahedral count drops by one, and the boundary geometry remains otherwise flexible, but now within a new combinatorial state.
A stronger transition occurs if deformation proceeds far enough that an apex–apex edge appears. In this regime the belt is no longer merely reclassified locally; it is shortened more substantially, and multiple tetrahedra are lost simultaneously. In an eight-vertex system, for example, this transition produces a loss of two or three tetrahedra (depending on how the polygon is decomposed) relative to the bipyramidal maximum. These changes are discrete and unavoidable: intermediate geometric states do not register combinatorially because the system skips them by reorganizing its incidence structure.
This behavior illustrates a general asymmetry between boundary and interior. Continuous geometric variation of the boundary can induce discrete changes in interior capacity, while fixing the combinatorial structure of the tetrahedral decomposition leaves substantial freedom in the realization of the boundary. The resulting “quantized” behavior of tetrahedral and boundary counts is not a physical effect, but a consequence of incidence-gated admissibility: counts remain constant under arbitrary rippling until a gate is crossed, at which point they jump.
5.8 Inferring external structure from internal decomposition
This subsection relates the internal tetrahedral decomposition of a genus–0 polyhedron to the combinatorics of its boundary surface. We recall the exact topological identities for tetrahedralizations of a 33–ball and their normal–form specialization (already established in Propositions 5.15.3), then record a useful incidence-based rule-of-thumb for reconstructing $(V,E,F)$ from interior decomposition data.
5.8.1 Exact Euler link and normal-form specialization
For any tetrahedral complex filling a 33–ball, the exact identity (Proposition 5.1)
$T-N_{i}+E_{i}-V_{i}=1$
always holds. In the normal form adopted throughout this manuscript, $V_{i}\equiv 0$ and every interior edge is an internal triangulation segment, so $E_{i}=S_{i}$ and hence (Proposition 5.3)